### v1

@@ Abstract
Artificial Intelligence (AI) has shown great promise in electrocardiogram (ECG) analysis and cardiovascular disease detection. However, developing a general AI-ECG model has been challenging due to inter-individual variability and the diversity of ECG diagnoses, limiting existing models to specific diagnostic tasks and datasets. Moreover, current AI-ECG models struggle to achieve comparable performance between single-lead and 12-lead ECGs, limiting the application of AI-ECG to portable and wearable ECG devices. To address these limitations, we introduce an ECG Foundation Model (ECGFounder), a general-purpose model that leverages real-world ECG annotations from cardiology experts to broaden the diagnostic capabilities of ECG analysis. ECGFounder is trained on over 10 million ECGs with 150 label categories from the Harvard-Emory ECG Database, enabling comprehensive cardiovascular disease diagnosis through ECG analysis. The model is designed to be both effective out-of-the-box and fine-tunable for downstream tasks, maximizing usability. More importantly, we extend its application to single-lead ECGs, enabling complex condition diagnoses and supporting various downstream tasks in mobile and remote monitoring scenarios. Experimental results demonstrate that ECGFounder achieves expert-level performance on internal validation sets for both 12-lead and single-lead ECGs, while also exhibiting strong classification performance and generalization across various diagnoses on external validation sets. When fine-tuned, ECGFounder outperforms baseline models in demographics detection, clinical event detection, and cross-modality cardiac rhythm diagnosis. ECGFounder not only enhances existing cardiovascular diagnostics capabilities but also facilitates integration into mobile and remote monitoring devices, enabling continuous, real-time cardiac health monitoring crucial for early detection and management of chronic heart conditions. The trained model and data will be publicly released upon publication through the Brain Data Science Platform (bdsp.io), promoting further research and development in this field. Our code is available at https://github.com/bdsp-core/ECGFounder.
@@ 1 Introduction
Electrocardiogram (ECG) is a crucial diagnostic tool that measures and records the electrical activity of the heart using electrodes placed on the skin.[2] ECG recordings are vital diagnosis and monitoring of cardiac health conditions.[14] In addition to common arrhythmias, ECGs can be used to diagnose conditions such as myocardial ischemia, myocardial hypertrophy, and myocardial infarction.[18] However, fully interpreting an ECG is complex and difficult. A typical ECG expert undergoes nearly 10 years of training, including medical school, internal medicine residency, and specialized ECG training.[2] Despite this, comprehensive and accurate interpretation of ECGs remains a challenge even for qualified experts. In real clinical scenarios an ECG may present multiple diagnoses simultaneously. For instance, a patient with coronary heart disease might also exhibit bundle branch block, ST-segment depression, and T-wave inversion. Even ECG experts may occasionally overlook or mislabel findings due to human error. There is an urgent need for an AI-powered automated model capable of identifying all ECG abnormalities.
Traditional ECG analysis focuses on extracting waveform and rhythm features from ECG signals for arrhythmia classification and ECG anomaly detection.[9, 28] These methods are simple and effective, but limited, as they rely on expert knowledge and prior experience, which may fail to capture complex patterns in the data. Manual feature selection can lead to missing important information or overfitting. To address this limitation, some studies have introduced signal analysis techniques, such as Fourier transform and wavelet transform, which allow for automatic feature extraction and classification of ECG signals. [32, 6] Compared to traditional ECG analysis, signal analysis methods offer higher accuracy. However, these models face challenges in handling and analyzing large-scale ECG data, and they can be computationally intensive.
Recent work has integrated deep learning models into ECG analysis, significantly improving diagnostic accuracy and efficiency.[42] Current deep learning approaches in ECG interpretation primarily employ Convolutional Neural Networks (CNNs)[16, 43, 25], Recurrent Neural Networks (RNNs)[34], Transformers[24] and their variants, which have proven effective in identifying common arrhythmias and conduction blocks.[10, 1, 30]
However, due to limitations of existing ECG databases in terms of sample size, patient numbers, and the variety of diagnoses, these models have not yet addressed the challenges of inter-individual variability and the diversity of diagnoses[23]. This restricts the models to performing only a limited set of diagnostic tasks and datasets. Additionally, current ECG models do not achieve the same level of performance for single-lead ECGs as they do for 12-lead ECGs.[17, 29] These challenges highlight the practical limitations of training an ECG model from scratch on small-scale datasets, making it difficult to extend the model to real-world ECG analysis. A promising approach is to leverage transfer learning and the strong generalization capabilities of foundation models to enhance the performance of ECG models.
@@
Foundation models have recently made significant strides in medical AI. In retinal disease diagnosis, the RETFound model, through pre-training on a large number of retinal images, has achieved excellent performance across various clinical diagnostic tasks.[41] In computational pathology, the UNI model was trained on a vast amount of whole-slide images, reaching expert-level performance in multiple cancer diagnostic tasks.[4] In these studies, foundation models are defined as large-scale AI models trained on extensive datasets, capable of adapting to a wide range of downstream tasks. Specifically, they meet the following criteria: 1) the pre-training dataset is large in scale, 2) the model has an enormous number of parameters, and 3) it can perform a wide range of downstream tasks. Although recent work has made substantial progress in many medical fields, there is currently no foundation model for ECGs.[36, 3]
@@
To fully leverage the advantages of foundation models, we develop an ECG Foundation Model—a groundbreaking approach aimed at revolutionizing ECG analysis. This foundation model, the first of its kind, is capable of diagnosing 150 types of cardiac abnormalities covered by existing ECG classifications by utilizing over ten million clinically annotated real-world ECG datasets. To address the inherent challenges of incomplete annotations in real-world data, we introduce a novel method for pre-processing and training on these annotations, ensuring robust performance even under sub-optimal conditions. Moreover, by training the single-lead ECG model based on lead augmentation, we are able to maintain high diagnostic performance on single-lead ECGs as well. We validate the model’s performance on both internal and external test sets, where it consistently matches expert-level diagnoses. In downstream task fine-tuning, we demonstrate ECGFounder’s versatility in addressing various tasks, including demographics detection, clinical event detection, and cross-modality diagnosis (Figure 1b). Specifically, we evaluate ECGFounder on 12 clinical tasks, such as ECG age regression and classification, sex detection, chronic kidney disease (CKD) detection, chronic heart disease (CHD) detection, regression and abnormal classification of left ventricular ejection fraction (LVEF), regression and abnormal classification of NT-proBNP, and atrial fibrillation detection based on photoplethysmography(PPG). The results of these downstream tasks highlight ECGFounder’s potential as a foundational model for the further development of AI-ECG models.
@@
Figure 1: ECGFounder is a general ECG encoder based on the RegNet architecture. (a) ECGFounder was trained on 10,771,552 ECGs with 150 types of ECG diagnostic labels. Due to the presence of missing diagnoses in real-world expert annotations, we implemented positive label augmentation by modifying the loss function of the pre-training method. For more details, see the Methods section. (b) ECGFounder was applied to 12 clinical downstream tasks, covering ECG diagnosis, demographics inference, clinical event detection, and cross-modality diagnosis. In comparison with baseline methods, ECGFounder achieved state-of-the-art performance across all tasks. (c) A comparison of the HEEDB dataset used by ECGFounder with other large ECG datasets. For more details, see Table LABEL:tb:data. (d) Frequency distribution of HEEDB labels. The x-axis is the number of samples in logarithm, and the y-axis is the label type. The label types are categorized based on their frequency into common, uncommon, and rare sets. The common set has a frequency greater than $10^{5}$. The uncommon set has a frequency greater than $10^{4}$ but less than $10^{5}$. The rare set has a frequency greater than $10^{3}$ but less than $10^{4}$. PPG., photoplethysmography; LVEF., left ventricular ejection fraction; CKD., chronic kidney disease; CHD., chronic heart disease.
@@ 2.1 Datasets and pre-processing
Our dataset, the Harvard-Emory ECG Database (HEEDB), is currently the largest open-access ECG dataset, containing 10,771,552 expert-annotated ECGs from 1,818,247 unique subjects.[19] These ECGs are predominantly 10-second, 12-lead clinical ECGs. The dataset includes annotations from cardiologists and ECG technicians paired with the ECGs. These annotations consist of discrete text reports that primarily describe the morphology, rhythm, and diagnostic information of the ECGs. Experts used the Marquette 12SL ECG Analysis Program (GE Healthcare) version 4 to assist with annotations,[12] which provides waveform parameters for doctors’ reference. The program offers many diagnostic categories, allowing doctors to simply click on corresponding category labels on the computer, avoiding manual entry of diagnostic statements. To extract descriptive information about ECGs from these discrete labels, we utilized regular expressions to parse the annotations, tallying each independent label. In total, there were 287 independent phrases. After reviewing with doctors, we removed phrases that were irrelevant to ECG descriptions while retaining meaningful phrases. Ultimately, we defined 150 meaningful labels, including rich diagnostic information as well as specific rhythm and morphological descriptions (See Supplement for more details about labels).
@@
Our external validation data comprised three large ECG databases (DB): The CODE-test DB [31], the PTB-XL DB[38], and PhysioNet Challenge-2017 DB [5]. The CODE-test DB is composed of ECG records from 827 patients across 811 municipalities in Brazil, collected by the Telehealth Network of Minas Gerais (TNMG). There are six common arrhythmia labels in this ECG DB, annotated by several experienced ECG experts.
@@
The PTB-XL dataset contains 21,837 clinical ECGs from 18,885 patients in Germany. Each ECG is a 10-second, 12-lead recording. The labels were reviewed and verified by two physicians. This dataset is currently one of the best publicly accessible ECG collections, both in terms of the number of samples and the quality of the labels.[38]
@@
The PhysioNet Challenge-2017 is a large single-lead ECG dataset. The ECG recordings were collected using the AliveCor device.[5] The training set includes 8,528 single-lead ECG recordings, with durations ranging from 9 to 60 seconds, while the test set contains 3,658 ECG recordings of similar lengths. The dataset requires classification of ECG recordings into normal rhythm, atrial fibrillation rhythm, other rhythms, and noise.
@@
Additionally, we utilized the MIMIC-IV-ECG DB to fine-tune our model for various downstream tasks. The MIMIC-IV-ECG dataset is part of the MIMIC series, focusing on the collection and analysis of ECG data.[7] MIMIC-IV-ECG originates from real clinical settings at Beth Israel Deaconess Medical Center (BIDMC) in Boston, USA, and contains a large number of clinical ECG recordings from patients treated in the intensive care unit (ICU). Moreover, ECG recordings in the dataset can be matched with the electronic health records of patients in the MIMIC-ED, allowing ECG data to be associated with specific conditions. Here, we linked several clinical downstream tasks, such as age, sex, chronic kidney disease (CKD), chronic heart disease (CHD), left ventricular ejection fraction (LVEF) and NT-proBNP with ECG data to explore the performance improvement of the fine-tuned ECGFounder model in detecting other diseases and clinical events. More details about the split and labels of MIMIC-IV-ECG DB can be found in Supplement.
@@
To explore ECGFounder’s generalization capabilities on other similar physiological signals, we fine-tuned and evaluated it using the DeepBeat dataset, a PPG-based atrial fibrillation (AF) detection dataset.[37] The dataset includes over 500,000 labeled signals from more than 100 individuals.
@@ 2.2 Model architecture
Considering these factors, we built our model architecture on top of the RegNet architecture.[15, 27] This structure begins with a stage-wise network design where each stage consists of a set number of blocks and channels that scale with network depth. This is beneficial for us to expand and design the blocks and channels of the ECG foundation model. Unlike traditional uniform scaling across layers, RegNet employs a quantized linear model to predict optimal widths and depths, ensuring efficient performance across a range of model sizes. The initial layers focus on capturing low-level features with fewer channels, which gradually increase as the network deepens, thus optimizing computational efficiency and model capacity. Following this, the model utilizes a series of bottleneck blocks that combine group convolutions and channel-wise attention mechanisms, enhancing the richness of representation in both temporal and spatial dimensions while controlling model complexity. This tailored configuration makes the model suitable for ECG data. More details about the model architecture can be found in the Supplement.
@@ 2.3 Training with real-world annotations: noisy, imbalanced, positive unlabeled
Unlike conventional ECG diagnostic models that typically use single-label classification methods, we employed multi-label classification during the training phase of the ECG foundation model. This approach aligns more closely with clinical practice, where an ECG diagnosis often consists of multiple diagnostic labels. For example, an abnormal ECG diagnosis might be something like normal sinus rhythm | premature ventricular complexes | premature ectopic complexes. Additionally, multi-label diagnostics better meet the application needs of clinicians. During training, utilizing a multi-label classification approach provides our model with rich semantic value and facilitates generalization to different annotation systems. However, the nature of multi-label annotations means that it is almost impossible to achieve perfect multi-label data, especially in cases like ECG where there are many diagnostic categories. Generally speaking, if a cardiology expert can annotate three classes simultaneously, it is considered excellent. However, the actual number of ECG diagnostic labels far exceeds three classes.[18] To address the challenge brought by incomplete ECG annotations, we introduce positive unlabeled (PU) learning method.
@@
PU learning is defined as when positive samples are present in the data but unlabeled, but the labeled positive samples are correct, meaning that the unlabeled samples are not necessarily negative examples.[40] Typically, PU learning leads to a severe imbalance in the predicted probability distribution: an ECG usually contains a few positive diagnostic labels, with the remaining labels being treated as negative, leading to a severe positive-negative imbalance where the negative samples far outnumber the positive samples for each label. The inherent scarcity of positive examples, combined with the missing labels, exacerbates the imbalance, making it more challenging to learn from the positive examples. When using conventional loss functions, such as the Binary Cross-Entropy (BCE) loss function, the model tends to learn from simpler samples, namely the true negative (TN) samples, while the more challenging samples, namely the false negative (FN) samples, are harder to fit. When the model is biased towards learning TN samples, its predicted probabilities tend to skew towards 0 rather than 1, which diminishes its ability to predict positive samples and weakens its potential to detect clinically important abnormalities.
@@
To enhance the model’s ability to fit representations of positive samples, we improved the multi-label classification loss function, enabling the model to correct missing labels by adjusting the weights of positive and negative samples.[39] Typically, for missing labels, a well-trained multi-label model’s predicted probability should be close to 1 rather than 0. Therefore, by applying a smaller weight to the loss of negative labels with predicted probabilities close to 1, we can mitigate the impact of missing labels. The loss function is given by:
@@
Our model training used AdamW to minimize the loss function, with an initial learning rate set to 0.001.[21] The learning rate decayed by a factor of 10 every 5 epochs. The trainable temperature parameter was initialized to 0. Training spanned a maximum of 20 epochs, with early stopping based on validation loss. We used a batch size of 1024.

### v2

@@ Abstract
BACKGROUND. Artificial intelligence (AI) has demonstrated significant potential in ECG analysis and cardiovascular disease assessment. Recently, foundation models have played a remarkable role in advancing medical AI, bringing benefits such as efficient disease diagnosis and cross-domain knowledge transfer. The development of an ECG foundation model holds the promise of elevating AI-ECG research to new heights. However, building such a model faces several challenges, including insufficient database sample sizes and inadequate generalization across multiple domains. Additionally, there is a notable performance gap between single-lead and multi-lead ECG analyses.
METHODS. We introduced an ECG Foundation Model (ECGFounder), a general-purpose model that leverages real-world ECG annotations from cardiology experts to broaden the diagnostic capabilities of ECG analysis. ECGFounder was trained on over 10 million ECGs with 150 label categories from the Harvard-Emory ECG Database, enabling comprehensive cardiovascular disease diagnosis through ECG analysis. The model is designed to be both an effective out-of-the-box solution, and a to be fine-tunable for downstream tasks, maximizing usability. Importantly, we extended its application to lower rank ECGs, and arbitrary single-lead ECGs in particular. ECGFounder is therefore applicable to supporting various downstream tasks in mobile and remote monitoring scenarios.
RESULTS. Experimental results demonstrate that ECGFounder achieves expert-level performance on internal validation sets, with AUROC exceeding 0.95 for eighty diagnoses. It also shows strong classification performance and generalization across various diagnoses on external validation sets. When fine-tuned, ECGFounder outperforms baseline models in demographic analysis, clinical event detection, and cross-modality cardiac rhythm diagnosis, surpassing baseline methods by 3 to 5 points in AUROC.
CONCLUSIONS. The ECG foundation model offers an effective solution, allowing it to generalize across a wide range of tasks. By enhancing existing cardiovascular diagnostics and supporting integration into mobile monitoring systems, it contributes to the advancement of the cardiovascular AI community and enables management of cardiac conditions.
@@ 1 Introduction
The Electrocardiogram (ECG) is a crucial diagnostic tool that measures and records the electrical activity of the heart using electrodes placed on the skin.1 ECG recordings are essential for diagnosis and monitoring of cardiac health conditions. However, fully interpreting an ECG is complex and requires significant training and time. A typical ECG expert undergoes nearly 10 years of training, including medical school, internal medicine residency, and specialized ECG training.2 In recent years, the advent of deep learning together with efforts to assemble relatively large databases of ECGs have seen some interesting progress in the field, extending ECG analysis beyond the traditional medical domains 3, 4, 5, 6. However, due to the lack of large-scale publicly available ECG databases with diverse diagnostic information, developing a general-purpose AI-ECG model remains a challenging task. Existing models are often confined to specific diagnostic tasks and datasets. This challenge highlights the practical limitations of training an ECG model from scratch on small-scale datasets, making it difficult to extend the model to real-world ECG analysis.
@@
Foundation models, with their strong generalization capabilities, provide a promising approach for enhancing the performance of AI-ECG models through transfer learning. Recently, these models have achieved significant advancements in the field of medical AI. In retinal disease diagnosis, the RETFound model, through pre-training on a large number of retinal images, has achieved excellent performance across various clinical diagnostic tasks.7 In computational pathology, the UNI model was trained on a vast amount of whole-slide images, reaching expert-level performance in multiple cancer diagnostic tasks.8 In these studies, foundation models are defined as large-scale AI models trained on extensive datasets, capable of adapting to a wide range of downstream tasks. Specifically, they meet the following criteria: 1) the pre-training dataset is large in scale, 2) the model has an enormous number of parameters, and 3) it can perform a wide range of downstream tasks.9
There have been several claims in the literature of developing foundation models for the ECG.10, 11 However, due to limitations of existing ECG databases in terms of sample size, patient numbers, the variety of diagnoses, and importantly, the demographics of the patients, these models have yet to address the challenges of diversity across regions, ethnic groups, and diagnostic variations.12 Moreover, to qualify as a foundation model, the trained model must be capable of generalizing to multiple domains outside of the initial training paradigm. 13 Additionally, current ECG models have significant performance degradation in single-lead ECGs compared to multi-lead ECGs.14
In this work, we propose the ECG Foundation Model, ECGFounder, which is capable of diagnosing 150 types of cardiac abnormalities, leveraging over ten million clinically annotated real-world ECGs spanning all existing ECG classifications. This represents the largest and most comprehensive ECG foundation model to date. Moreover, we demonstrate how the model is applicable to a wide range of tasks in varying domains, a fundamental requirement for foundation models.We provide the medical machine learning community with an accessible ECG foundation model that offers an effective out-of-the-box solution and fine-tuning capabilities. Compared to traditional AI-ECG models trained from scratch, ECGFounder offers a novel approach that achieves superior performance through fine-tuning. This advancement has the potential to drive the future development of AI-ECG technology.
@@
To address the inherent challenges of incomplete annotations in real-world data, we introduce a novel method for pre-processing and training on these annotations, ensuring robust performance even under sub-optimal conditions. Moreover, by training the single-lead ECG model based on lead augmentation, we are able to maintain high diagnostic performance on single-lead ECGs as well. We validate the model’s performance on both internal and external test sets, where it consistently matches expert-level diagnoses. In downstream task fine-tuning, we demonstrate ECGFounder’s versatility in addressing various tasks, including demographics detection, clinical event detection, and cross-modality diagnosis (Figure 1b). Specifically, we evaluate ECGFounder on 12 clinical tasks, such as ECG age regression and classification, sex detection, chronic kidney disease (CKD) detection, chronic heart disease (CHD) detection, regression and abnormal classification of left ventricular ejection fraction (LVEF), regression and abnormal classification of NT-proBNP, and atrial fibrillation detection based on photoplethysmography (PPG). The results of these downstream tasks highlight ECGFounder’s potential as a foundational model for the further development of AI-ECG models.
The trained model and labeled data will be publicly released upon publication through the Brain Data Science Platform (bdsp.io), huggingface and PhysioNet, promoting further research and development in this field.
@@
Figure 1: ECGFounder is a general ECG encoder based on the RegNet architecture. (a) ECGFounder was trained on 10,771,552 ECGs with 150 types of ECG diagnostic labels. Due to the presence of missing diagnoses in real-world expert annotations, we implemented positive label augmentation by modifying the loss function of the pre-training method. For more details, see the Methods section. (b) ECGFounder was applied to 12 clinical downstream tasks, covering ECG diagnosis, demographics inference, clinical event detection, and cross-modality diagnosis. In comparison with baseline methods, ECGFounder achieved state-of-the-art performance across all tasks. (c) A comparison of the HEEDB dataset used by ECGFounder with other large ECG datasets. (d) Data used in model development and validation for the ECG diagnosis and downstream tasks. PPG., photoplethysmography; LVEF., left ventricular ejection fraction; CKD., chronic kidney disease; CHD., chronic heart disease.
@@ 2.1 Datasets and pre-processing
Our dataset, the Harvard-Emory ECG Database (HEEDB), is currently the largest open-access ECG dataset, containing 10,771,552 expert-annotated ECGs from 1,818,247 unique subjects.15 These ECGs are predominantly 10-second, 12-lead clinical ECGs. The dataset includes annotations from cardiologists and ECG technicians paired with the ECGs. These annotations consist of discrete text reports that primarily describe the morphology, rhythm, and diagnostic information of the ECGs. Experts used the Marquette 12SL ECG Analysis Program (GE Healthcare) version 4 to assist with annotations, which provides waveform parameters for doctors’ reference.16 The program offers many diagnostic categories, allowing doctors to simply click on corresponding category labels on the computer, avoiding manual entry of diagnostic statements. To extract descriptive information about ECGs from these discrete labels, we utilized regular expressions to parse the annotations, tallying each independent label. In total, there were 287 independent phrases. After reviewing with doctors, we removed phrases that were irrelevant to ECG descriptions while retaining meaningful phrases. Ultimately, we defined 150 meaningful labels, including rich diagnostic information as well as specific rhythm and morphological descriptions (See Supplement for more details about labels).
@@
Our external validation data comprised three large ECG databases (DB): The CODE-test DB 17, the PTB-XL DB 18, and PhysioNet Challenge-2017 DB 19. The CODE-test DB is composed of ECG records from 827 patients across 811 municipalities in Brazil, collected by the Telehealth Network of Minas Gerais (TNMG). There are six common arrhythmia labels in this ECG DB, annotated by several experienced ECG experts.
@@
The PTB-XL dataset contains 21,837 clinical ECGs from 18,885 patients in Germany. Each ECG is a 10-second, 12-lead recording. The labels were reviewed and verified by two physicians. This dataset is currently one of the best publicly accessible ECG collections, both in terms of the number of samples and the quality of the labels.18
@@
The PhysioNet Challenge-2017 is a large single-lead ECG dataset. The ECG recordings were collected using the AliveCor device.19 The training set includes 8,528 single-lead ECG recordings, with durations ranging from 9 to 60 seconds, while the test set contains 3,658 ECG recordings of similar lengths. The dataset requires classification of ECG recordings into normal rhythm, atrial fibrillation rhythm, other rhythms, and noise.
@@
Additionally, we utilized the MIMIC-IV-ECG DB to fine-tune our model for various downstream tasks. The MIMIC-IV-ECG dataset is part of the MIMIC series, focusing on the collection and analysis of ECG data.20, 21 MIMIC-IV-ECG originates from real clinical settings at Beth Israel Deaconess Medical Center (BIDMC) in Boston, USA, and contains 800,035 clinical ECGs from 161,352 patients treated in the intensive care unit (ICU). Moreover, ECG recordings in the dataset can be matched with the electronic health records of patients in the MIMIC-ED, allowing ECG data to be associated with specific conditions. Here, we linked several clinical downstream tasks, such as age, sex, chronic kidney disease (CKD), chronic heart disease (CHD), left ventricular ejection fraction (LVEF) and NT-proBNP with ECG data to explore the performance improvement of the fine-tuned ECGFounder model in detecting other diseases and clinical events. More details about the split and labels of MIMIC-IV-ECG DB can be found in Supplement.
@@
To explore ECGFounder’s generalization capabilities on other similar physiological signals, we fine-tuned and evaluated it using the DeepBeat dataset, a PPG-based atrial fibrillation (AF) detection dataset.22 The dataset includes over 500,000 labeled signals from more than 100 individuals.
@@ 2.2 Model architecture
Considering these factors, we built our model architecture on top of the RegNet architecture.23, 24 This structure begins with a stage-wise network design where each stage consists of a set number of blocks and channels that scale with network depth. This is beneficial for us to expand and design the blocks and channels of the ECG foundation model. Unlike traditional uniform scaling across layers, RegNet employs a quantized linear model to predict optimal widths and depths, ensuring efficient performance across a range of model sizes. The initial layers focus on capturing low-level features with fewer channels, which gradually increase as the network deepens, thus optimizing computational efficiency and model capacity. Following this, the model utilizes a series of bottleneck blocks that combine group convolutions and channel-wise attention mechanisms, enhancing the richness of representation in both temporal and spatial dimensions while controlling model complexity. This tailored configuration makes the model suitable for ECG data. More details about the model architecture can be found in the Supplement.
@@ 2.3 Training with real-world annotations: noisy, imbalanced, positive unlabeled
Unlike conventional ECG diagnostic models that typically use single-label classification methods, we employed multi-label classification during the training phase of the ECG foundation model. This approach aligns more closely with clinical practice, where an ECG diagnosis often consists of multiple diagnostic labels. For example, an abnormal ECG diagnosis might be something like normal sinus rhythm | premature ventricular complexes | premature ectopic complexes. Additionally, multi-label diagnostics better meet the application needs of clinicians. During training, utilizing a multi-label classification approach provides our model with rich semantic value and facilitates generalization to different annotation systems. However, the nature of multi-label annotations means that it is almost impossible to achieve perfect multi-label data, especially in cases like ECG where there are many diagnostic categories. Generally speaking, if a cardiology expert can annotate three classes simultaneously, it is considered excellent. However, the actual number of ECG diagnostic labels far exceeds three classes.25 To address the challenge brought by incomplete ECG annotations, we introduce positive unlabeled (PU) learning method.
@@
PU learning is defined as when positive samples are present in the data but unlabeled, but the labeled positive samples are correct, meaning that the unlabeled samples are not necessarily negative examples.26 Typically, PU learning leads to a severe imbalance in the predicted probability distribution: an ECG usually contains a few positive diagnostic labels, with the remaining labels being treated as negative, leading to a severe positive-negative imbalance where the negative samples far outnumber the positive samples for each label. The inherent scarcity of positive examples, combined with the missing labels, exacerbates the imbalance, making it more challenging to learn from the positive examples. When using conventional loss functions, such as the Binary Cross-Entropy (BCE) loss function, the model tends to learn from simpler samples, namely the true negative (TN) samples, while the more challenging samples, namely the false negative (FN) samples, are harder to fit. When the model is biased towards learning TN samples, its predicted probabilities tend to skew towards 0 rather than 1, which diminishes its ability to predict positive samples and weakens its potential to detect clinically important abnormalities.
@@
To enhance the model’s ability to fit representations of positive samples, we improved the multi-label classification loss function, enabling the model to correct missing labels by adjusting the weights of positive and negative samples. Typically, for missing labels, a well-trained multi-label model’s predicted probability should be close to 1 rather than 0. Therefore, by applying a smaller weight to the loss of negative labels with predicted probabilities close to 1, we can mitigate the impact of missing labels. The loss function is given by:
@@
Our model training used AdamW to minimize the loss function, with an initial learning rate set to 0.001. The learning rate decayed by a factor of 10 every 5 epochs. The trainable temperature parameter was initialized to 0. Training spanned a maximum of 20 epochs, with early stopping based on validation loss. We used a batch size of 1024.